Two interior angles of a polygon are 88˚ and 99˚, and each of the rest equals to 163˚. Find the number of sides of the polygon.
Answer by ikleyn(52776)

.

88 + 99 + (n-2)*163 = (n-2)*180


88 + 99 + 163n - 2*163 = 180n - 2*180


88 + 99 - 2*163 + 2*180 = 180n - 163n


221 = 17n


  n = 221/17 = 13.


ANSWER.  13 sides.

Solved.